OpenStudy (anonymous):

PLLEZZ HELP!! describe and correct the error given the functions f(x)=3x-5 and g(x)=1/3-4/3x 15. f(5)=3(5)-5 5f=15-5 5f=10 f=2

OpenStudy (tkhunny):

"f(5)" is a functional structure. There is nothing to factor, here. "f" doesn't mean anything, just sitting there by itself.

OpenStudy (anonymous):

I have to correct the error in question 15

OpenStudy (anonymous):

You can correct it by reading the above statement?

OpenStudy (tkhunny):

Without looking at the error example, how would you work the problem: If f(x) = 3x - 5, what is f(5)?

OpenStudy (anonymous):

f(5) = 3(5)-5 f(5) = 15-5 f(5) = 10.

OpenStudy (anonymous):

is this correct?

OpenStudy (tkhunny):

Done. You corrected the errors.

OpenStudy (anonymous):

so is wat i put right or wrong?

OpenStudy (tkhunny):

Your task was, "correct the error". I said, "You corrected the errors." I'm not sure what more I can say. Move on to the next problem.

OpenStudy (anonymous):

sorry lmao. thank u tho